Mosaic_Engineering_Status_10_13_2008_thru_10_19_2008 by yaoyufang

VIEWS: 5 PAGES: 15

									From: Stein, Jack
Sent: Sunday, October 26, 2008 4:08 PM
To: Johnson, Bob; Smelser, Ron
Cc: Gay, Earl; Iyanar, Lokeshvar; Mehta, Sushant; Mori, Jayesh; Palaniswami, Siddharth; Tikekar, Rewa;
Venkataraman, Anuradha; Castaneda, Armando; Desgraves, Sean; DeWitt, Charles; Dyer, Rodney;
Edgecombe, Jason; Rhodes, Terry; Simineau, Wes; Whisler, Dave
Subject: Mosaic Computing Status for week ending Oct 13,2008 – Oct 19,2008

Mosaic Computing Status for week ending Oct 13,2008 – Oct 19,2008
Jack Stein:
 Key highlights
     A Mosaic Windows update was done to include:
               Microsoft Office & Windows XP patches
               MSN Live Messenger
               Thunderbird 2.0
     Prepared presentation on PCS for AIM. Attended AIM meeting Thur afternoon thru Sunday.
     We continue to address the problem with use of the Centra software and OneNote in the Woodward 140
        Lab. A concerted effort is in place to work with Dr. Ivan Howitt and Alan Hege to get the issues
        addressed. We are trying to replicate the problem in other podiums, get more memory into the system,
        instrument the system in Woodward 140 and monitor the network. We are making some headway on all
        these fronts.
     Working with Jay Dominic get twice per year PC buy/replacement process (instead of once per year). We
        have about 50 Mosaic Windows primary workstations that need replaced (from 2005).
     A non-Mosaic Windows PC is set up in CARC 235 connected to the network and running logged in to
        Groove. This enables all members of Dr. Gary Wilson’s class to upload and share documents.
     Continued to work on the weather station on the CARC roof and the acquisition of data from it so it can be
        displayed on our CARC 101 digital display. This is an initiative from Dr. Saunders of Civil.
     The effort by facilities to determine the cause and the repair of the CARC window water leakage has been
        awaiting resolution (mainly funding).
     We found out that ITS did not really get the Wireless Access Point (WAP) set up in the Smith 274B (Tony
        Brizendine’s office area) area as they indicate. After ITS did install the WAP we determined that
        Wireless Access is operational and the service is now in place.
     We plan to shut down the Centra EMeeting and discard all our licenses as FcTel is not interested in the
         transfer all of our Centra licenses to them.
     We continue dealing with very large number of application install for the faculty. Some are not really
        appropriate for College wide deployment.
     Our testing effort with NCSU’s VCL has been successful and the first effort with installation of the
        IMPLAN software for Dr. Ertunga Ozelkan and a few of his graduate students has been completed. So far
        this has worked out well.
     Terry is continuing with Primary and Secondary workstation deployment for faculty and staff.
     Continue to monitor and track the plethora of projects across Mosaic, PCS and with Support.
     We are working to evaluate VMWare and the ESX module as the kernel (replacing Linux) to determine
        how an RWS system performs with many more Mosaic XP guests on the same box we currently only can
        support 2 guests using VMWare Server. In parallel, we are received a loaner Dell server (r805 with ESX
        in firmware) using iSCSI disk array for the purpose of determining how many virtual Windows XP guest
        images can be supported by this configuration. Benchmark testing by our students has been completed.
        Analysis of the data, fine tuning the parameters and a re-test is in progress.
     Work is in progress to build a Red Hat Linux 64 bit release by this Dec 2008.
     Preparing for the audit visit – set for November 5,2008
     Dan Rowe has met with me and he is requesting a dash board view of all workstation, server rooms, storage
        rooms as part of an audit of how much is for research and how much is for academic use. I’ll be working on
        this.
     Attended the ITS Recharge meeting. Our mission is complete and a report will be issued. Recommendation
      is to have ITS charges removed from the current F & A Rate and to launch a separate task force on a
      strategy for ITS recharge.

  Customer Support:
   Armando Castaneda and Sean Desgraves continues working with the team on support issues.
   Armando is continuing to work on a project to monitor departmental and research Mosaic labs on a
      different SLA using Nagios, Lab Patrol and a web cam. The SEEM lab in Colvard 2046 is being
      monitored on a pilot basis. A SLA was drafted for SEEM and we are monitoring this departmental lab.
   Armando continues developing and shaping the Mosaic Help Desk and Mosaic TA staff. TA are writing
      How To pages on products.
   Mosaic Incidents Tracking for the week (these are ones logged in our incident tracking system, we handle
      many other incidents that are not logged and are resolved quickly):
                Total Incidents = 1665 Open/Waiting=72 Closed=1593

Infrastructure and Services:
        The disk farm service was announced on August 25. The disk farm will provide COE faculty
           researchers a place to store large data sets (200GB – 2 TB), with back up support, on our AFS Mosaic
           file system. Dave Whisler build the data base for this service. Howie Fang is using the service. No one
           else has come forth requesting this service. I plan to announce this again and see if there is any more
           users.
        Terry and I are meeting with the High Capacity Work Group headed by Chuck Price. At this point the
           COE is the only place where large disk space with back up is available as a service. Other groups are
           trying to get ITS to provide the same service. Chuck Price is preparing a report on this initiative.
        The 5 R300 Dell Servers are in. 3 test machines for Jason / TS. 1 Machine for VMWare
           VirtualCenter. 1 Machine for TS test LAMP / Wiki. We are planning on displacing the 2 745’s in the
           CARC server room once these are installed. The 745’s can be used to refresh the Mosaic TS machines
           in CARC

Personal Computer Support:
      o I met with Roberta Knapp of Apple to review the price increases that Apple triggered on Oct 1 when
           the special Apple laptop promotion ended in September. We also discussed that any Apple
           announcements that would impact our offering need to be discussed with me. New Apple pricing
            flyer and other items will be delivered. We are also working with Apple on a holiday laptop
           promotion.
      o I met with Mike Kelly of Lenovo to review the status of the laptop initiative. Mike will be working
           with me on a holiday promotion.
      o I was supposed to meet with Dell but it was cancelled and we are trying to reschedule.
      o Continued effort to work on the buildup of the Personal Computing Support group.
      o We have been working with Dave Whisler to make enhancements to the Work Order data base for the
           SOC initiative. Also working with shop inventory system to handle loaner laptop and projectors.
      o The phone number for the Personal Computer Support is x73355. Office location is in Smith 202a.
      o Armando Castaneda, Ian Pugsley, Sean Desgraves and I are meeting regularly and reviewing the SOC
           project and initiatives.
      o Dell Statement of Work ( SOW) arrived Jesh from legal gave me his feedback.
                o PCS Work Orders Summary Report
                                                PCS Work Orders Summary Report
                         WO's opened between                     07/01/08    and          10/26/08
                         Opened by Month
                             July 2008                                                     16
                                                           % Total Opened              13.33%
                             August 2008                                                   23
                                                           % Total Opened              19.17%
                               September 2008                                              52
                                                            % Total Opened            43.33%
                               October 2008                                               29
                                                            % Total Opened            24.17%
                                                             Total Opened                 120
                          Last Worked by Month
                               July 2008                                                   14
                                                        % Total Last Worked          11.67%
                               August 2008                                                 21
                                                        % Total Last Worked          17.50%
                               September 2008                                              56
                                                        % Total Last Worked          46.67%
                               October 2008                                                29
                                                        % Total Last Worked          24.17%
                                                          Total Last Worked               120
                          Closed by Month
                               July 2008                                                  14
                                                            % Total Closed           12.07%
                               August 2008                                               21
                                                            % Total Closed           18.10%
                               September 2008                                            56
                                                            % Total Closed           48.28%
                               October 2008                                               25
                                                            % Total Closed           21.55%
                                                             Total Closed                116


 Digital Signage:
             I am currently working with Mike Hermann and am managing the insertion of the contents in the
             Nexus driven Players.
          The Digital Signage with Nexus is set up in Hawthorne. It appears that I do have access the
             Hawthorne Mac Mini via vnc. Working with Cathy Blat on contents and access to the building for me.
          Working with Mike Herman to compile contents for ECE to use on the Digital Signage.
          An initiative is underway to have all UNCC based contents available for the digital signs.
          Nexus is pushing to have us move from our current software to the NCS software.
          I am training Rewa and Lokesh when time permits on the Nexus product.


Classroom Lecture Capture and Distance Education:
         I am working to support Tony Brizendine and Steve Clark on the Distance Education project for ET in
            Smith 202. We have made good progress with the PanOpto Coursecast 2.0 software. We have started
            to do testing with one of the video cameras we ordered; we expect the second camera to be available
            week of Oct 20 and we will test with 3 cameras (Canon zr900_. We are also testing integration with
            Blackboard. I provide Tony with weekly updates. I held a presentation to ET faculty on the status of
            the project. Steve and I are scheduling a status meeting for Oct 29 for a more widespread group of
            people to determine what is the campus position on making the Panopto lecture capture go from testing
            into production. I met with Steve Clark and Lee to review various topics on Panopto. We also had a
            conference call with Panopto to review some of our understanding and misunderstanding about the
            product..
            Working on SEEM Lecture Capture initiative. They are interested in Panopto for later when it is
             ready.
            Harish Cherukuri will migrate to PanOpto when it is ready.

 WEB Projects:
    After much grief with the User Interface of the MEES web site we were able to have a compromise of what
      the UI should look like. The compromise was to take what Anu and Sid came up with, have a graphic
      designer that was hires by UNCC help clean it up. Sid an Anu will be working to create this for ME.
    The Mosaic Web site is being re-assessed and we will rework the menu and contents to make it more
      intuitive and easier to find things.
    We plan to start testing another wiki open source product as the xwiki is not working out for us in our
      extensive testing efforts on the Nexus project.
    Working with Larry Sharp to set up ASME as a virtual web site.
    Project is in progress to re-do the web form for appeal by suspended students (for Karina Franco). It was
      released for Karina to test and assess.
    Work is in progress to update LiveHelp web application (we are about 2 years behind). It was released but
      later on we had to pull it back due to issues with the MySQL server.
    Work is in progress to update the Web calendar application. It was tested and is ready for release.
    Work is in progress to evaluate use of YouTube for placement of our videos
    Work is in place to have our COE web news be available as RSS feed.
    Work has started to create a web based application to manage our PCS inventory (i.e. loaner laptops,
      projectors, etc), our consumable (toner, paper, etc), guest accounts, and vacation planner.
    ITS (Carter Heath) has blocked the release of IP addresses for our 2 virtual host systems. He is requesting
      that we update some of our software components on the COE Web but we will plan to do that in Dec
      2008. I am working with Jason on getting us new SSL certificates for the 2 secured COE web sites:
      https://www.coe.uncc.edu https://secure.coe.uncc.edu when we get these and Jason has these in place on
      our COE web server I will announce to all of COE. Right now when accessing these sites we are
      recommending to press continue when getting a warning message.
      Hit Counts on our existing Virtual Tour for COE: Since May 4,2007: 3630

Miscl:
        We continue to use Nagios open source web based software to monitor all our services in real time. The TA
         are also using this tool to monitor any issues with our services.

New Initiatives:
              We are researching the Wake-on-Lan feature that will let us turn on Dell Windows PC systems
                 that are turned off. Smith 260 is configured this way. With this feature we can now go “green” by
                 shutting down our Mosaic workstations (remotely) in labs that are closed and then turning them on
                 when we open a lab.
              Wes worked with Steve Clark’s group to deploy an elegant method of logging off a user who
                 forgot to do so in a podium. Deployed in Woodward 140 and Alan Hege has no issues with it.
                 Deployed in Smith 21,218,219 podiums and Help Desk to test. We plan to release this to all
                 Mosaic based podiums.
              The virtual Windows XP system has been released as Beta to Brian Anderson and Harish
                 Cherukuri as well as Ron Smelser. A User Guide showing the installing and managing of this
                 offering has been produced. So far Harish and Ron have been very enthusiastic with this product.
                 Harish has found some performance issues with AFS access and we have diagnosed the issue and
                 provided a fix that appears to have fixed the issue. We are working to get a release update so we
                 can seed Tony Brizendine and Steve Patterson.
              The Kodak digital picture frame (10”) with wireless support is installed in Smith 202a (PCS
                 Help Desk) We received and are testing the picture frame with direct Ethernet support (as well as
                 wireless) from Dlink. We plan to test this to see if we can use it for Mosaic labs/classrooms to
                 show room schedule and other content.
              PDA strategy for the COE. We seeded Dean Johnson with the Apple Ipod Touch as a PDA.
                 Dean Johnson is reporting initial success. We distributed the 4 additional iPod PDAs (1 for Ron
                   Smelser, 1 for Tony Brizendine. 1 for Dr. Chung-Suk Cho and 1 for Martin Kane). There is
                   already a line of eager faculty and staff who would like to get an iPod PDA. 6 more Apple iPod
                   Touch PDA have arrived and have been or will be distributed to: Bill Heybruck, Yogi Kakad,
                   Helene Hilger, Bev Guessford, Jack Stein, Patty Tolley.
                  Looking at a different methodology for event calendar on COE Web.
                  Looking at a different methodology of how room reservation are managed. Looking at wireless
                   digital photo displays mounted on wall or using of Exchange shared calendar for the handling the
                   rooms (this is how HSS and C of Ed does it). Also looking at R25 as another possibility.
                  We will be testing the Linux OS as the file server OS – replacing over a 2-3 year period the Sun
                   OS based file servers.
                  Wes has been migrating the license services off ls-sm1 to ls-me1

Applications:
    We plan to purchase Adobe CS-4 in Nov 2008 and install it on Mosaic.
    We are working on the Abaqus 6.8 on Linux for Harish
    The Comsol RF module has been released for Marco Scipioni.
    We plan to remove Eudora from Mosaic.
    Working on the Prolog software for Dr. Chung-Suk Cho of ET. We are running into difficulties with this
         package.
    ME has decided to fund the renewal of Adam software and we are waiting for the paper work.
    Working with Dr. John Gresser on PCB design software that is not in the Cadence suite but may be with
         Multisim
    Released Implan software and data on VCL for Ertunga Ozelkan. So far no issues.
    Working on the software install for Dr. Mesbah Uddin: STAR CCM+
    Working to get software for Dr. Mesbah Uddin: CFD-ACE
    We have re-installed the Synopsys software on Unix and Linux. ECE will find an alternative product
         solution for their research and teaching needs. Work is in place with Ron Smelser and ECE to get sign off
         to remove splash page.
    No activity on the Aveva software. Rodney has been looking for cycles to get up to speed on this
         configuration. Areva has not been responding to any request for assistance. Rodney and I met with Ian
         Piggott of Aveva to see if he can help us kick start the PDMS installation. The ball is really in the
         departments to decide the classes and syllabus that incorporates this product.


Software purchase, renewal status:

           Expiration      Software                    Costs                Notes

                  6/1/08   GRL Weap                              Contacted Brian Anderson if it needs renewed; he needs
                                                                 to let us know
                 7/31/08   Respondus                             Chad needs to get updated version from FcTel
                 8/31/08   Algor                                 Need to check with Dr. Chen – he renews this.
                  9/6/08   ADAMS                        $1500    Working with Dr. Cuttino
                 12/1/08   Labview                      $2500    Being handled by Chuck Price

Planned for Nov 2008 Update:
        Microsoft Windows patches and Office patches

In the queue
          Windows Update Agent (prerequisite)
         Firefox 3.0
         PRS Clicker Software
         ABAQUS 6.8
          SmartBoard
          TNT Identity Client Removal
          JInitiator Removal (multiple versions)
          Origin 8.4
          ArcGIS 8.3 Removal (need to begin communicating this packages removal to faculty)
           AFS Client / Kerberos Client (assuming testing goes well)
           Oracle 11g – In Queue
           Visual ModFLOW Pro 4.3 – Low Priority
          P-Spice / OrCAD / Layout suite replacement
          Prolog
          Maple 12 on all 3 platforms
          PanOpto Coursecast Recorder and Editor software for lecture capture


On Hold
          Acrobat Pro
           Timberline Office – Current version is functioning postponing until possibly next summer
           Hyperworks 9 – Pushing out until either Winter break or next summer.
                            License issue driving update has been resolved.

We are considering using the ITS Citrix service for packages like SAS/JMP as well as ArcGIS.




Chad DeWitt:
   -   Prepared iPod Touches for deployment
           o Unpackaged iPods and connected to iTunes for initialization
           o Recorded MAC, user machine, serial number and purchase date into DevDB
   -   Worked with Wes on License Server Maintenance (LS-SM2, LS-ME1, LS-ME2, LS-ME3)
           o LS-SM2
                     removed several deprecated processes
                     Standardized on FlexLM license server options
                     replaced bad hard drive
                     Installed several patches
                     Created ghost image
                     Installed new license server maintenance script for log file management
                     created micro archives of license and admin directories.
                     Updated respective documentation
                     Tested applications after server restart to ensure license server functionality.
                     cleaned up XFER, backup, and dept/MOSAIC/ls-sm2 folders
           o LS-ME1
                     Installed several patches
                     Standardized on FlexLM license server options
                     Copied VMWare image files to AFS
                     created micro archives of license and admin directories.
                     Tested applications after server restart to ensure license server functionality.
                     cleaned up XFER, backup, and dept/MOSAIC/ls-me1 folders
                     compressed and archived previous backup
           o LS-ME2
                     Installed several patches
                     Copied VMWare image files to AFS
                     created micro archives of license and admin directories.
                     Tested applications after server restart to ensure license server functionality.
                     cleaned up XFER, backup, and dept/MOSAIC/ls-me2 folders
                     compressed and archived previous backup
              o  LS-ME3
                       Installed several patches
                       Standardized on FlexLM license server options
                       Created ghost image
                       created micro archives of license and admin directories.
                       Tested applications after server restart to ensure license server functionality.
                       cleaned up XFER, backup, and dept/MOSAIC/ls-me3 folders
                       replaced USB Hub
             o MOSQL
                       Installed several patches
                       Created ghost image
    -    Worked on Assessment of ProLog software with Wes
             o After complications we did get the ProLog license server running on our license server
             o Prolog is very poorly designed and requires RW access for all users to the server where license is
                 hosted.
             o DB Portion is also poorly designed requiring install of optional software on MS-SQL server
    -    Looked at Chris Maynard’s PC with Wes
             o reconfigured Keyboard and mouse to attempt to alleviate issue
    -    Worked with Rodney to “demote” ADCSM1
             o Moved FSMO roles from ADCSM1 to ADCSM3
             o Verified proper DC operation
             o Added necessary IP address for ADCSM3 to AD DNS
             o Moved GC from ADCSM1 to ADCSM3
    -    Worked with Wes on Prolog License Manager
    -    Worked on Dr. Raquet’s printer issue with Armando and Wes
             o Checked connectivity
    -    Worked with Rodney and Wes on Update
             o Assembled update
             o Tested update
    -    Got software from InterlinkElectronics what was missing from ePad device
    -    Packaged ePad-Ink driver

PTO
   -     December 22, 23
   -     January 2

Wes Simineau:
---------------------
Worked on maintaining / backing up Windows servers with Chad
            LS-SM2
                      removed several deprecated processes
                      Standardized on FlexLM license server options
                      replaced bad hard drive
                      Installed several patches
                      Created ghost image
                      Installed new license server maintenance script for log file management
                      created micro archives of license and admin directories.
                      Updated respective documentation
                      Tested applications after server restart to ensure license server functionality.
                      cleaned up XFER, backup, and dept/MOSAIC/ls-sm2 folders
            LS-ME1
                      Installed several patches
                      Standardized on FlexLM license server options
                      Copied VMWare image files to AFS
                      created micro archives of license and admin directories.
                      Tested applications after server restart to ensure license server functionality.
                    cleaned up XFER, backup, and dept/MOSAIC/ls-me1 folders
                    compressed and archived previous backup
          LS-ME2
                    Installed several patches
                    Copied VMWare image files to AFS
                    created micro archives of license and admin directories.
                    Tested applications after server restart to ensure license server functionality.
                    cleaned up XFER, backup, and dept/MOSAIC/ls-me2 folders
                    compressed and archived previous backup
          LS-ME3
                    Installed several patches
                    Standardized on FlexLM license server options
                    Created ghost image
                    created micro archives of license and admin directories.
                    Tested applications after server restart to ensure license server functionality.
                    cleaned up XFER, backup, and dept/MOSAIC/ls-me3 folders
                    replaced USB Hub
          MOSQL
                    Installed several patches
                    Created ghost image
Worked on Assessment of ProLog software with Chad
          After complications we did get the ProLog license server running on our license server
          Prolog is very poorly designed and requires RW access for all users to server where the license is hosted.
          DB Portion is also poorly designed requiring install of optional software on MS-SQL server
Looked at Chris Maynard’s PC with Chad
          reconfigured Keyboard and mouse to attempt to alleviate issue
Followed up with Dr. Raquet on Outlook issue
Installed COMSOL RF Module on 4 workstations for Marco in Grigg
Discussed digital signage project with Jack
Proposed and Reviewed AppDB modifications with Dave/Jack/Sid for handling VCL & Citrix Application
tracking.
Began investigating package for Cadence and StarCCM+
Following up on ProE issue in Duke 242
Worked with Rodney and Chad to modify the baseline / test update
Misc Updates to AppDB.
Installed VMWare Player for PCS group on two workstations per Ian’s request
Worked with Chad / Armando on Duke Printer in 323
Worked with TAs on Groove machine setup

PTO
   -     December 22, 23
   -     January 2


Jason Edgecombe:
----------------------
Done:
* shutdown matlab license daemon on ls-sm1
* fixed a missing library on Linux for Dr. Adams
* helped Deb Sharer with WebPE
* removed legalese banner from Synopsis
* worked with Steve Stoy form ITS on Symantec issue
* resolved a support issue where the Mosaic menu wasn't appearing for a user on Solaris
* followed up with Dr. Mukherjee. He doesn't need a surplus machine.
* tweaked nss_db script to make the output quieter and send fewer emails.
Waiting:
* Waiting on Dr. Mukherjee for Wind River info.
* Dean Tolley will use a WebWork class for Fall 2008 that is hosted by the Math Dept. We will revisit in the Dec
2008 to see if we need to have our own server.

In Progress:
* working on SSL name error on www.coe SSL web sites with two-level names - blocked by ITS security until we
upgrade apache+php
* as part of 64bit Linux testing, test each Mosaic application on 64 bit and make an upgrade strategy.
* Started working Redhat 5.2 64-bit build -- got a box to build with 64bit RHEL5U2.
* helping Chris McDaniels with Webwork courses
* get R200 build with Linux for Wes -- will try 64bit. - requires drivers in RHEL5U1 or higher.
* AFS testing with outlook and OST files - installed latest client build
* investigate dropped packets from afs fileserver under load -- happens w/Outlook OST
* Earl: Installed Comsol (32/64) on 64bit Linux. -- waiting on testing
* nss_db causes gnome login to take too long on 64bit linux. Incident filed with RedHat.
* Earl installed on Solaris: ProENGINEER WildFire 4 -- waiting on testing.
* Earl installed Abaqus on 64bit Linux -- needs testing
* Earl installed Star CCM+ on 32bit Linux for Dr. Uddin - waiting on testing.
* working Dr. Soliman to help him access the ~saesb web page for editing.

In queue:
* Nagios work - needs extra custom plugins
* Linux as a file server project.
* upgrade php and apache on web server.
* reorganize/remodel shared ME folders for Robin Coger -- after Fall starts
* install on Linux/Solaris: ADAMS
* upgrade thunderbird on Solaris
* upgrade thunderbird on linux
* troubleshoot why 32bit Maple 10 doesn't work on 64bit Linux.
* fix DNS server issues -- upgrade or remove
* add AFS rxdebug "calls waiting" metric to nagios
* fix apache logs archiving script to handle new ssl config
* shutdown samwrap license daemons on ls-sm1
* upgrade opnet on solaris/linux
* upgrade comsol on solaris
* install a better PDF reader on solaris - requested by Wes
* point cadence & pspice on solaris/linux to new license server
* tweak alert timeouts in nagios for MosaicAnywhere machines to 2 hours. Add cron job to alert every 6 while in
maintenance.
* install Cadence on Linux
* install CCM+ on 64bit Linux for Dr. Uddin
* install CFD-ACE on Linux for Dr. Uddin
* setup www.asme.uncc.edu vhost on coe-web by end of Fall 2008 for Larry Sharpe.

Postponed until Summer 2009 or later:
* RT Work - a test install is running. Working on customization.


Dave Whisler:
-------------------
Mosaic:
                o     Attended weekly staff meeting
                o     Created accounts for new TA’s in XWiki and put them in the MOSAIC and PCS groups.
                o     Granted Oracle roles PCS TA, MOSAIC TA, and INV TA to new Mosaic TA’s.
                o     Worked with Jack to improve his Powerpoint Presentation for the AIM meeting this weekend.
SupportDB:
             o   Looked into errors reported by Sean. Appears database is working as it should based on the
                 screenshots he sent.
             o   Created new report (PC Registrations Summary and Detail reports) requested by Sean Desgraves
             o   Need to write documentation of Oracle back-end and Access front-end
             o   Wrote some SQL queries to pull WorkOrder statistics for Jack’s presentation at the AIM meeting

XWiki:
             o   Reviewed release notes for 1.6 (new version). Brand-new GUI, but needs time to mature.
                 Holding off on upgrade until later release.

AppDB:
                Met with Wes and Jack to discuss modifications to accommodate new needs (tracking of apps on
                 VCL and Citrix systems).
                Added new fields to APP_PACKAGE table: PCS_AVAIL_VCL_FLAG,
                 PCS_AVAIL_CITRIX_FLAG, ON_VCL_FLAG, ON_CITRIX_FLAG, PCS_LICENSE_NOTES
                Recompiled triggers and views referencing APP_PACKAGE table
                Re-linked APP_PACKAGE table link in MS-Access front-end
                Added all of the new fields to the App_Package Form in MS-Access front end
                Added new logic to the CREATE_APP_FILES Oracle Procedure to handle these new fields and
                 dump them to the various Mosaic/PCS webfiles.
                Added new logic to the Help Files Upload VBA code on MS-Access front end.

AdminDB:
             o   Met with Toby Rufty.
             o   Working on a long list of enhancements to existing reports.
             o   Working on combining a couple of separate reports into one report using run-time parameters to
                 make it more flexible.
             o   Will add a new table and several new fields to the database.
             o   Will modify the front-end forms to handle the new table and new fields

MySQL:
                Upgraded my installation of MySQL Workbench

Oracle:
                Read chapters 1-4 of Oracle PL/SQL Programming book
                Upgraded my installation of SQL Developer
                Assessing new Oracle CPU (Critical Patch Update) for October.

MRS DB:
             o   Looking into pulling info from Banner on all COE courses (name, number, section, instructor,
                 schedule, building, room)
             o   Ran Course Worksheet in Report Central out to excel spreadsheet for Jack
             o   Did research on capabilities of R25.
             o   Had good conversations with Donna McLain and Lucian Wilhelm on its abilities for our use
             o   Had Lucian remove access to R25 from former COE staff.

Shop Inventory DB:
         Gave demo to Sid and Anu, explained functionality and workflow of the system

DeviceDB:
        Made some improvements for Rodney.

PCS:
             Need to write documentation of MySQL back-end

MRS:
             Need to write documentation of MySQL back-end


Terry Rhodes:

•          Vacation Monday thru Thursday
•          Recovered Sun machine and monitor from Woodward 467
•          Installed standalone machine Woodward 467 for Gloria Elliot
•          Picked up old PC and monitor from Woodward 467 to surplus
•          Picked up machine from Sara Watson for testing and evaluation
•          Removed Mosaic PC from Smith 245 and hooked up digital sender to the
           network
•          Created Applications volumes
•          Restored volume for cdoss4 to look for missing file
•          Working to perform normal job duties
•          Performed weekly AFS backups to tape
•          Working to reorganize storage rooms

Vacation           Nov 24 - 26
                    Dec 22 & 23
                    Jan 2, 2009

Rodney Dyer:
-----------------
week ending 10/19/08:

* Help for Dr. Raphael Tsu.
  Getting a very big PowerPoint Presentation via email.

* Help for Dr. Shen Guo Wang.
  ArcGIS data format problem from NCDOT?

* Fix of "prcsinfo.vbs" script bug that prevented the
  dataprobe script from killing "hung" ssh processes.

* Mosaic XP Maintenance Update.

* Download, scripting, and testing of security patches for
  Office 2007 and Windows XP.

* Added special defrag (watchdog) monitoring code into
  the update script to prevent extremely long defrags on
  some machines.

* Continued OneNote testing for Dr. Howitt.

* Helped Steve Stoy with gathering data for Symantec to
  find problem with Endpoint Protection.

* Purchase order from Jeffrey Altman.

* Continued work with Chad to migrate old ADCSM1 server
  to new ADCSM3 Active Directory rack server.
* Continued testing of IMAP problems with Dan Dombchewskyj.



Armando Castaneda:

.        Provided Karina Franco with floor by floor Smith building layout
.        Opened Woodward labs
.        Staff meeting
.        Update lab schedules
.        Update Mosaic tickets
.        Close Mosaic tickets
.        Laserprint Company came out and updated the firmware on the 249 LaserJet printer
.        Assigned 5 training account to MAPS office for use on October 25 th
.        Followed up with Liz Scott and Chris Maynard concerning helpdesk assistance incident
.        Opened Magic ticket for Jeff Raquet concerning VLAN problem
.        Order more paper
.        Informed Dan Rowe of carpet in Smith 202 not allowing the door to close
.        Canceled Nextel service
.        Worked with Charis Mayhorn from ITS on Duke 323 printer problem (no solution yet)

Sean Desgraves:
Weekly report for the week of: 10/13/08
•          Went to Staff meeting on Tuesday
•          Reviewed Camera Playback of 24 hour labs for non-business hours
•          Took various shifts in the PCS helpdesk to monitor and assist with customer issues and requests
•          Closed various Mosaic/PCS Tickets/WOs
•          Received from Mike Herman the order of 500 copies of the updated PCS brochure
•          Delivered 400 of the updated PCS to Karina Franco as requested
•          Setup a VMWare Windows XP image for testing
•          Assisted Armando in testing Centra with Outlook for the podium in Woodward 140
•          Returned banners and tablecloth to Lenovo through Team World Wide (shipping)
•          Received 6 new iPod Touches from Chad to deliver to various people
•          Delivered iPod Touch to Bill Heybruck
•          Delivered iPod Touch to Yogendra Kakad
•          Delivered iPod Touch to Jack Stein
•          Provided printer research for Drs. Daneshvar and Tranjan as requested for use with future laptop
purchases
•          Scheduled delivery of the iPod Touch for Bev Guessford and Helene Hilger for next week
•          Reserved for Amy Kitts 4 laptops and a printer to be checked out Friday of next week
•          Followed up with Ron Smelser for assistance with specs for a MacBook Pro purchase
•          Tested Tony Brizendine’s office for network connectivity
•          Provided Drs. Daneshvar and Tranjan with a quote from Apple for a 15” MacBook Pro
•          Contacted Robin Coger to schedule a time to pick up her laptop to install campus software
•          Received 2 Tablet PCs from ITS for Sharon Green/Churlzu Lim – contacted them with a list of software
to be installed
•          Contacted Patty Tolley to schedule delivery of her iPod Touch


Siddharth Palaniswami:

Attended COE web meeting
Met Carrie Whitehead regarding room reservations
Met with MEES and Maggie to discuss the MEES layouts
Designed new layouts for MEES
Working on making the graphic layout into HTML
Working on writing an aggregator script for aggregating probedata to generate reports on software usage.


Anuradha Venkataraman:
I was working on the following last week:

Worked on virtual tours
Met with MEES on redesign
Worked on new layouts for MEES
Worked on converting graphic to HTML
Worked on software list to generate new help pages.


Weekly Status report Sushant Mehta:
-----------------------------------------------------------------------------
Activities for week ending Oct 13,2008-Oct 19,2008

Summary:

1)Completed documentation for web calendar.
2)Completed RSS feeds for edit and delete functionality.
3)Started documentation for RSS feeds.
4)Completed ticket creation for Mosaic account activation.
5)Tested and deployed updated Live Help.


Weekly report for of Jayesh Mori:

1> Completed Karina's Suspension Appeal App
2> Deployed Groove Server and Tested.
3> Deployed Groove Workspace of Dory and tested
4> Updated Live Help
5> Done WebCalender final changes
6> Helped Sushant on RSS Feeds.


Earl Gay:

Done:
- Added OpenOffice to 64-bit Linux install.
- Added all 32 bit EPEL repo packages to the 64 bit repo, and upgraded those packages to their 64 bit versions.
- Downloaded and added i386 Java to mosaic extra repo; documented procedure to enable said java in Firefox.
- Wrote script to determine differences in actual user disk quota and the quota listed in the database.
- Sent output from quotacompare script to Mr. Whisler to allow him to sync the database with the actual quotas.
- Added a2ps to 64 bit build.
- Went with Mr. Edgecombe to troubleshoot a library issue in Woodward.

Progress:
- Worked on generating a list of packages needed to be migrated to 64bit install.
- Installed 64-bit Opnet 14 on Linux. (Waiting on testing)
- Installed Pro|E Wildfire 4.0 (Waiting on testing)
- Installed StarCCM+ (Waiting on testing)
- Installed Cadence IC5. (Waiting on testing)
- Worked on debugging dbus: tried different architecture versions, and tweaking configurations, etc.
- Looked into dbus source code for possible error with hardcoding passwd information.
- Started copying volumes for Opnet, StarCCM+, and Wildfire4 onto their permanent volumes.

Lokeshvar Iyanar:
* I have wrapped the editing process of the video that we had captured the other day on the presentation. I had made
inferences on the scenario on how to make offline recordings and had uploaded the same over the network later. It
worked just fine. I tried logging off when the recorder is uploading a video.

* I tried to do an edit the 20 minute lecture that I captured on Tuesday and while it is encoding disconnect the PC
from the network. reconnect to the network 30 minutes later. Goal is to see if encoding completed while recorder
PC is disconnected

* I had recorded a 10 minute video, audio, desktop and power point slides to a memory stick or to an external USB
drive. Then I took the USB drive or memory stick to another recording PC and point to this drive and had the
lectures uploaded from that.

* I had made tests on how to download lectures onto ipods and had made inferences on that. I had also tested the
RSS feed of the lectures that we had captured.

* I had gone through the entire user manual of both the Canon ZR900 and the tripod to understand more about its
features and how to make the most out of the two. I had tested the cameras while having them mounted them on the
tripods.

Timesheet for the last week

Mon                2:10 PM                    5:20 PM
Tue               12:00 PM                    6:00 PM
Thur              12:20 PM                    7:45 PM
Fri                2:30 PM                    6:00 PM




Rewa Tikekar:
My weekly status report is as follows:

Monday: 12:30 to 6:30PM
Worked on modifying the documents.

Tuesday: 1:00PM to 5:30 PM
1. Meeting conveying the functionality and status of Panopto.
2. Performing some tests for panopto regarding upload when Log off and recording when not on network.

Thursday: 11:30 AM to 7:00PM
Performed the tests mentioned in the schedule for 16th Oct.
1. To edit the 20 minute lecture and while it is encoding, disconnect the PC from the network. reconnect to the
network 30 minutes later. Goal is to see if encoding completed while recorder PC is disconnected (we were told
this would be ok).
2. During a 5 minutes recording disconnect the camera. The act of disconnecting the camera should end the
recording with an error.
3. Record a 10 minute video, audio, desktop and power point slides to a memory stick or to an external USB drive.
Then take the USB drive or memory stick to another recording PC and point to this drive (same drive letter and path
as was done on original recording PC) and see if it will upload (we were told this would be ok). Be sure the external
drive is USB 2.0 and is de fragmented.
Also carried out some tests with ipod.

Friday: 2:00 PM to 6:30PM

Read the entire user guide for the Canon zr900. We mounted the camera onto the tripod. We were also able to
connect the camera to the Apple Micro Mini and view it on the Panopto Recorder.



 Jack M. Stein
Assistant Dean and Director of Engineering Computing The University of North Carolina at Charlotte
226 Cameron Research Building
9201 University City Boulevard
Charlotte, NC 28223
jmstein@uncc.edu
704-687-6625
412-848-6534 (cell)
704-687-2352 (fax)

								
To top